OpenStudy (anonymous):

check 25-[2+5x-3(x+2)]=-3(2x-5)-[5(x-1)-3x+3]

OpenStudy (anonymous):

\[25-(5 x-3 (x+2)+2)=-3 (2 x-5)-(5 (x-1)-3 x+3) \]\[29-2 x=17-8 x \]\[x=-2 \]